# Simulating Human-like Daily Activities with Desire-driven Autonomy

### v2

@@ C.2 Indoor environment
Appendix D Case Study for action diversity
@@
Appendix E Mapping for Calculating expected value
@@
Appendix F Algorithm to calculate expected value
@@
To ensure that the language model outputs content related to ”desire” as consistently as possible, our numerical values are all integers. Since our descent rate is a floating-point number, we have designed a probability-based descent. The probability in each step can be calculated as follows :
@@
Appendix G Value Description
@@
Appendix H Prompt for Qualitative Value Description
@@
Appendix I Prompt for Value Update
I.1 Value Update Question
@@ I.3 Reflection Question
Appendix J Component designs in D2A
J.1 Goal setting for D2A
@@
J.2 Prompt for Activity proposal
@@
J.3 Prompt for Activity evaluation
@@
J.4 Prompt for Activity selection
@@
Appendix K Habits generation prompt for baselines
@@ 8. Eating Habits:
Appendix L Prompt used in babyAGI
L.1 Initialization prompt
@@
L.3 activity creation prompt
@@
Appendix M Prompt used in LLMob
M.1 Creation of activity that likely to do
@@
M.2 Motivation Generation Prompt
@@
Appendix N Prompt used in ReAct
@@
Appendix O Experiment conducted with Qwen model
@@
To evaluate the adaptability of our framework across different large language models, we tested it with Qwen 2.5 72B (Yang et al., 2024), which can also be integrated through the Ollama interface. As demonstrated in Figure 14, our framework, when running on the Qwen 2.5 72B base model, successfully provides desire-driven action guidance for the agent.
Appendix P An Example of D2A action sequences for indoor environment
@@
Appendix Q An Example of ReAct action sequences for indoor environment
@@
Appendix R An Example of BabyAGI action sequences for indoor environment
@@
Appendix S An Example of LLMob action sequences for indoor environment
@@
Appendix T An Example of Human action sequences for indoor environment
@@
Appendix U Raw action sequences of the indoor environment
U.1 D2A
@@
Appendix V The Examples of preprocessed action sequences for the outdoor environment
V.1 D2A
@@ V.4 LLMob
Appendix W Evaluation Prompt
W.1 GPT4o
@@
Appendix X Indoor Environment With Randomness
X.1 Pre-defined item lists
@@
X.2 Environment Generating Prompt
@@
X.3 Results
Figure 15: The win rates among 4 agents in randomly initialized indoor environments. Each point represents the win rate of the agent on the vertical axis when compared to the agent on the horizontal axis. The diagonal values are set to 0.5.

### v3

@@ C.2 Indoor environment
Appendix D Case Study for in Outdoor Environment
@@
D.3 The Win Rate Heatmap in Outdoor Environment
Figure 13: The win rates among 4 agents in outdoor environments. Each point represents the win rate of the agent on the vertical axis when compared to the agent on the horizontal axis. The diagonal values are set to 0.5.
As shown in Figure 13, our agent framework also outperforms the three baseline agents in the outdoor environment.
Appendix E Indoor
Appendix F Mapping for Calculating expected value
@@
Appendix G Algorithm to calculate expected value
@@
To ensure that the language model outputs content related to “desire” as consistently as possible, our numerical values are all integers. Since our descent rate is a floating-point number, we have designed a probability-based descent. The probability in each step can be calculated as follows :
@@
Appendix H Value Description
@@
Appendix I Prompt for Qualitative Value Description
@@
Appendix J Prompt for Value Update
J.1 Value Update Question
@@
Appendix K Component designs in D2A
K.1 Goal setting for D2A
@@
K.2 Prompt for Activity proposal
@@
K.3 Prompt for Activity evaluation
@@
K.4 Prompt for Activity selection
@@
Appendix L Habits generation prompt for baselines
@@ 8. Eating Habits:
Appendix M Prompt used in babyAGI
M.1 Initialization prompt
@@
M.3 activity creation prompt
@@
Appendix N Prompt used in LLMob
N.1 Creation of activity that likely to do
@@
N.2 Motivation Generation Prompt
@@
Appendix O Prompt used in ReAct
@@
Appendix P Experiment conducted with Qwen model
@@
To evaluate the adaptability of our framework across different large language models, we tested it with Qwen 2.5:72B (Yang et al., 2024), which can also be integrated through the Ollama interface. As demonstrated in Figure 15, our framework, when running on the Qwen 2.5:72B base model, successfully provides desire-driven action guidance for the agent.
Appendix Q An Example of D2A action sequences for indoor environment
@@
Appendix R An Example of ReAct action sequences for indoor environment
@@
Appendix S An Example of BabyAGI action sequences for indoor environment
@@
Appendix T An Example of LLMob action sequences for indoor environment
@@
Appendix U An Example of Human action sequences for indoor environment
@@
Appendix V Raw action sequences of the indoor environment
V.1 D2A
@@ V.5 Human
Appendix W The Examples of preprocessed action sequences for the outdoor environment
W.1 D2A
@@
Appendix X Evaluation Prompt
X.1 GPT4o
@@
Appendix Y Indoor Environment With Randomness
Y.1 Pre-defined item lists
@@
Y.2 Environment Generating Prompt
@@
Y.3 Results